OpenStudy (anonymous):

What is (are) the soulution(s) of the equation? x square minus 7x = -12

OpenStudy (anonymous):

\[x^2-7x=12\]So, subtract 12 from both sides to get,\[x^2-7x+12=0\]I don't know what methods you use to solve (quadratic, completing the square, factoring), but factoring here I think is simplest. To factors that multiply to give +12 and add to give -7 are -3 and -4. So\[(x-3)(x-4)=0\]This is true when \[x=3, x=4\]